I have 51 handle bars and 116 wheels. Using all the parts how many tricycles and bicycles can I assemble?
Answer by stanbon(75887)

I have 51 handle bars and 116 wheels. Using all the parts how many tricycles and bicycles can I assemble?
-----------------------------------------
Let # of tricycles be "t".
Let # of bicycles by "b".
---------------------------
Handle bar equation: t + b = 51
Wheel equation.....: 3t + 2b = 116
-------------------------------------
Rewrite the equations:
3t + 3b = 153
3t + 2b = 116
------------------
Subtract 2nd from 1st to solve for "b":
b = 37 (# of bicycles)
------
Substitute into t+b = 51 to solve for "t"
t + 37 = 51
t = 14 (# of tricycles)
===========================
Cheers,
Stan H.